OpenStudy (anonymous):

PLEASE HELP!!!: Two sides of a rectangle have lengths 20 and 32. Find the ratio of the perimeter to the longer side 1)13:4 2)4:13 3)5:8 4)8:5

jimthompson5910 (jim_thompson5910):

Perimeter: P = 2L + 2W P = 2*32+2*20 P = 64+40 P = 104 So the ratio of the perimeter to the longer side is 104:32, but this simplifies because we can divide each component by the GCF 8 to get the final answer of 13:4